if (3,3) lies on line joining points (h,0) and (0,k) then:
  1. h+k=9
  2. 1/h+1/k=1/3
  3. hk=3
  4. 3h-3k=1

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Slope of the line joining (h,0) and (0,k) = -k/h.
Line joining (h,0) and (0,k) is given by:
y – k = - kx/h => kx + hy = kh.
Now we know that this line passes through the point (3,3). Hence the equation of the line should satisfy (x,y) = (3,3).
Therefore we have 3k + 3h = kh.
On diving this expression with kh, we arrive at 3/h + 3/k = 1.
This can be rearranged as 1/h + 1/k = 1/3.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)